We will discuss the following reflector problem. Given a parallel light beam emitted from a bounded domain in a hyperplane of n+1 dimensional Euclidean space and a receiver surface as well as prescribed in advance intensities for the source domain and the receiver, can we design a reflector surface that redistributes the emitted light accordingly?